OpenStudy (mikeymaximum):

What is –3 2/3· (–2 1/4)? A.–8 1/4 B.–6 1/6 C.8 1/4 D.8 3/4

OpenStudy (ajspeller):

Convert those mixed numbers into improper fractions. Then multiply the numerators and denominators and simplify if possible.

OpenStudy (hockeychick23):

-11/3*-9/4= 99/12= 8 3/12= 8 1/4